OpenStudy (anonymous):

Simplify (x^k + 1)^2(–x^k)^3 (3x). A. –3x^5k + 2 B. –3x^5k + 3 C. 3x^5k + 2 D. 3x^k + 7

jhonyy9 (jhonyy9):

inside first parentheses just k is exponent ?

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (anonymous):

Yesss

OpenStudy (anonymous):

no k+1 is exponent

OpenStudy (anonymous):

\[(x ^{k+1}) ^{2} (-x ^{k})^{3} (3x)\]

jhonyy9 (jhonyy9):

(x^k + 1)^2(–x^k)^3 (3x).= =(x^(2k +2)(-x^3k)*(3x) = = x^2k *x^2 *(-x^3k)*(3x) = = x^(2k+2) * (-3x^(3k+1) = = - 3x^(2k+2+3k+1) = = - 3x^(5k+3)

jhonyy9 (jhonyy9):

so B. is right ok ?

OpenStudy (anonymous):

Yessss I understand!!! thanks a lot!